Derive the momentum of inertia component of a rotating rigid body

1
Expert's answer
2022-01-14T09:48:14-0500

Assume we have a point mass m rotating at distance r from the axis of rotation at speed v. Then, by definition, the moment of inertia is


"I=\dfrac L\omega=\dfrac {[mv·r]}{[v/r]}=mr^2."